# Benchmarking Chinese Commonsense Reasoning of LLMs: From Chinese-Specifics to Reasoning-Memorization Correlations

### v1

@@ 4.2 Prompt Strategies
Translate-EN: We use the DeepL api [^12] https://www.deepl.com/translator to translate our benchmark into English, and then use English CoT for reasoning Zhang et al. (2023a).
XLT: The template prompt Huang et al. (2023a) is used to change the original question into an English request, solve it step by step, and finally format the answer for output .
The examples for each prompt strategy are in Figure 8 in Appendix D. For all prompt strategies, we use the 3-shot setting.
@@ 5.1 Integrated Reasoning Performance
We show the performance of the 19 LLMs on CHARM reasoning tasks in Table 4. We only choose one representative prompt strategy: XLT for English LLMs and ZH-CoT for Chinese-oriented LLMs, which is based on our empirical conclusion in §5.2. The LLMs’ performance on the 7 aspects of the Chinese commonsense domain are shown in Table 10 in Appendix F.
Commonsense Domain We found that the LLMs exhibit inconsistent performance in the global and Chinese commonsense domains. The rankings of the English LLMs dropped in the Chinese domain compared to the global domain. For instance, GPT4 ranks first in the global domain, but in the Chinese domain, Qwen-72B outperforms all, pushing GPT4 to the second. In the Chinese domain, the performance of LLaMA-2-70B is even worse than many Chinese-oriented LLMs in the 6B-7B size range.
@@ 5.2 Prompt Strategy Selection
We tested the combinations of the 19 LLMs and the 5 prompt strategies in CHARM reasoning tasks. Detailed results are in Table 8 and Table 9 in Appendix E. To draw some empirical conclusions, we analyze along the following two dimensions:
@@
We average the $19\times 5$ LLM-prompt combinations along the above two dimensions, and the obtained results are in the Table 5. From the LLM dimension, it’s clear that various LLMs prefer different prompt strategies: XLT consistently excels for English LLMs among the 5 strategies, while for Chinese-oriented LLMs, despite some complexity, ZH-CoT generally performs best. From the commonsense domain dimension, strategies that use English for reasoning (like XLT, Translate-EN, etc.) are suitable for the global domain; however, ZH-CoT generally performs better in the Chinese domain.
@@ 5.3 Integrated Reasoning vs Memorization
We evaluated the correlation between the integrated reasoning and the memorization on the MRI tasks, as mentioned in §3.2. The average performance of the LLMs on the 4 MRI tasks is in Figure 3. Detailed performance on each task is in Figure 9 in Appendix G.1.
@@
•  Type II: High memorization and medium integrated reasoning ability. GPT3.5 and all Chinese-oriented LLMs below 30B belong to this type. It’s worth noting that some LLMs have high memorization performance, but relatively poor integrated reasoning ability.
•  Type III: Ultra-high memorization and high integrated reasoning ability. This category includes GPT4 and the three Chinese-oriented LLMs that exceed a size of 30B.
@@
In addition, we also evaluated the correlation between memorization and integrated reasoning during the LLM pre-training process, details can be found in Figure 10 in Appendix G.2.
@@ 16 LLaMA-2-13B LLaMA-2-70B (↓1) LLaMA-2-70B (↓1)
Table 6: Leaderboard on the MRI tasks. We propose two methods, i.e. FRMM and MIC, to compare the LLMs’ memorization-independent reasoning, as detailed in Appendix H. The arrows and numbers in brackets in the last two columns indicate changes in ranking order relative to the second column.
We propose two methods to compare the LLMs’ memorization-independent reasoning on the MRI tasks. The detail of the Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM) is in Appendix H.1, and Memorization-Independent Battles among LLMs (MIB) in Appendix H.2. The results are in Table 6. When comparing the leaderboards for integrated and memorization-independent reasoning, Type III LLMs rank at the forefront and the Type I rank at the end in all leaderboards. There is a slight variation in the ranking order within the three types of LLMs.
For the in-depth analysis, we chose Vicuna-13B, Qwen-7B, Qwen-72B as representatives for Type I, II, and III LLMs, and filtered out the reasoning questions in the MRI tasks, only keeping those with correct answers to the related memorization questions, same as the FRMM in Appendix H.1. This ensures the LLM has sufficiently memorized the commonsense knowledge required for the retained reasoning questions, thereby minimizing the impact of memorization on reasoning. There are totally 500 reasoning questions in the 4 MRI tasks, and the numbers of the retained are 106, 323 and 402 for Vicuna-13B, Qwen-7B and Qwen-72B respectively, as shown in Table 7.
Error Types If the LLMs provided incorrect answers for the retained reasoning questions, these errors can be referred to as memorization-independent reasoning errors. We conducted the manual review and analysis of their reasoning process, and classified the errors into 4 main categories.
@@
We show examples of each type of errors in Figure 12 in Appendix I. The distribution of these error types are shown in Figure 4.
@@
Vicuna-13B-v1.5
Type I
500 106 54
Qwen-7B
Type II
500 323 117
Qwen-72B
Type III
500 402 63
@@ 6 Conclusion
This paper introduces CHARM, the first benchmark designed to comprehensively and thoroughly evaluate LLMs’ commonsense reasoning in Chinese. CHARM encompasses two counterpart commonsense domains, global and Chinese-specific, with the carefully selected tasks. We evaluated the representative prompt strategies for improving LLMs’ reasoning ability, and the empirical findings significantly enhances and supplements the conclusions of previous studies. CHARM comprises closely-interconnected reasoning and memorization tasks, helping to reveal the intrinsic correlation between memorization and reasoning of LLMs. We have evaluated the strengths and weaknesses of different LLMs and conducted the detailed analysis of memorization-independent reasoning abilities. We hope that CHARM’s approach to studying the correlations between memoriztion and reasoning can serve as a reference for similar research in other fields.
@@
The automation of the construction and evaluation of CHARM needs further improvement, including the following: (1) Most of the questions in CHARM Chinese domain are manually constructed by the author. This limits the number of benchmark questions and the range of knowledge pieces covered. (2) Regarding memorization-independent reasoning, we chose only 3 LLMs as representative and manually categorized the types of errors within CHARM. In future research, we could employ robust LLMs, like GPT4, for automated error classification and statistical analysis.
@@
This work involves human annotation. We have provided appropriate compensation for all annotators. The total cost of annotation for the project is about 2.2k RMB. For all annotators, we explicitly inform them about the use of the data and require them to ensure that the questions included in CHARM do not involve any social bias, ethical issues or privacy concerns during the annotation process.
@@ Appendix C Question Examples of the Memorization Task in CHARM
Appendix D Examples of Prompt Strategies
@@
Appendix E Detailed Evaluation Results of 19 LLMs with 5 Prompt Strageties on Reasoning Tasks
@@
Appendix F Performance of LLMs on Chinese Commonsense Knowledge Aspects
@@
Appendix G Correlation of Memorization and Integrated Reasoning
G.1 Detailed Correlations of Memorization and Integrated Reasoning on the 4 MRI Tasks
@@
G.2 Correlation of Memorization and Integrated Reasoning throughout the LLM pretraining
@@
With the increase in the number of tokens during the training process, the model’s memorization ability quickly reaches a high level (in fact, there is no particularly obvious difference between the results of the first checkpoint and the final results). This is because the knowledge involved in our task setting is the most basic commonsense, and thus widely and abundantly exists in various Chinese training corpora.
@@
Appendix H Leaderboard of Memorization-Independent Reasoning
It is non-trivial to acquire and compare the memorization-independent reasoning abilities of the LLMs. Intuitively, we can filter the reasoning questions by only retaining those whose related memorization questions are all correctedly answered by every LLMs. This approach ensures that each LLM has memorized the commonsense knowledge necessary for the retained reasoning questions. However, when we apply this process to all the 19 LLMs, only 28 reasoning questions remain out of the original 500 in the MRI tasks, which is obviously insufficient in number and lacks diversity, thereby introducing a high degree of uncertainty due to randomness.
Therefore, we propose two slightly more complex methods, one called Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM), the other is Memorization-Independent Battles among LLMs (MIB).
H.1 Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM)
This method is relatively simple, but has some flaws to a certain extent. For each LLM, we select reasoning questions based on its performance in memorization tasks: only retaining reasoning questions for which all related memorization questions are answered correctly. It’s clear that, after individual filtration, different LLMs will retain different reasoning questions, and even differ in the number of retained reasoning questions, as shown in the “# retained” column in Table 11.
Then, we calculate the accuracy of the retained reasoning questions for each LLM, and the results are shown in the “Retained Acc” column in Table 11. The LLMs are then ranked based on the accuracy, producing the leaderboard shown in the penultimate column of Table 6.
@@ H.1 Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM)
To overcome this, we have proposed the MIB method.
H.2 Memorization-Independent Battles among LLMs (MIB)
To overcome the shortcomings of the FRMM method, we refer to the pairwise battle method adopted in LLM evaluation Zheng et al. (2023). By tallying each LLM’s performance in a “round-robin” tournament of pairwise match-ups, we rank the performance of the LLMs.
Specifically, we select two LLMs at a time and filter the MRI task’s reasoning questions based on the performance of these two LLMs in memorization tasks. We only retain the reasoning questions whose related memorization questions are correctly answered by both LLMs. In this way, the two LLMs are battled under fair conditions. We then calculate the accuracy of these two LLMs on the retained reasoning questions separately, and compute the difference in accuracy as the battle score between the two models.
As shown in Figure 11, the element $E_{ij}$ represents the accuracy of LLM $i$ minus the accuracy of LLM $j$ during the battle between the two LLMs. For a total of 19 LLMs, we average each model’s scores from the 18 battles they participated in as their final score, as shown in Table 12.
Finally, we rank the LLMs based on these scores to produce the leaderboard shown in the last column of Table 6.
Appendix I Memorization-Independent Reasoning Errors

### v2

@@ 4.2 Prompt Strategies
Translate-EN: We used the DeepL api [^14] https://www.deepl.com/translator to translate our benchmark into English, and then used English CoT for reasoning Zhang et al. (2023a).
XLT: The template prompt Huang et al. (2023a) was used to change the original question into an English request, solve it step by step, and finally format the answer for output .
The examples for each prompt strategy are in Figure 8 in Appendix E. For all prompt strategies, we use the 3-shot setting.
@@ 5.1 Integrated Reasoning Performance
We show the performance of the 19 LLMs on CHARM reasoning tasks in Table 5. We only choose one representative prompt strategy: XLT for English LLMs and ZH-CoT for Chinese-oriented LLMs, which is based on our empirical conclusion in §5.2. The LLMs’ performance on the 7 aspects of the Chinese commonsense domain are shown in Table 11 in Appendix G.
Commonsense Domain We found that the LLMs exhibit inconsistent performance in the global and Chinese commonsense domains. The rankings of the English LLMs dropped in the Chinese domain compared to the global domain. For instance, GPT-4 ranks first in the global domain, but in the Chinese domain, Qwen-72B outperforms all, pushing GPT-4 to the second. In the Chinese domain, the performance of LLaMA-2-70B is even worse than many Chinese-oriented LLMs in the 6B-7B size range.
@@ 5.2 Prompt Strategy Selection
We tested the combinations of the 19 LLMs and the 5 prompt strategies in CHARM reasoning tasks. Detailed results are in Table 9 and Table 10 in Appendix F. To draw some empirical conclusions, we analyzed along the following two dimensions:
@@
We averaged the $19\times 5$ LLM-prompt combinations along the above two dimensions, and the obtained results are in the Table 6. From the LLM dimension, it’s clear that various LLMs prefer different prompt strategies: XLT consistently excels for English LLMs among the 5 strategies, while for Chinese-oriented LLMs, despite some complexity, ZH-CoT generally performs best. From the commonsense domain dimension, strategies that use English for reasoning (like XLT, Translate-EN, etc.) are suitable for the global domain; however, ZH-CoT generally performs better in the Chinese domain.
@@ 5.3 Integrated Reasoning vs Memorization
We evaluated the correlation between the integrated reasoning and the memorization on the MRI tasks, as mentioned in §3.2. The average performance of the LLMs on the 4 MRI tasks is in Figure 3. Detailed performance on each task is in Figure 9 in Appendix H.1.
@@
•  Type II: High memorization and medium integrated reasoning ability. GPT-3.5 and all Chinese-oriented LLMs below 30B belong to this type. It’s worth noting that some LLMs have high memorization performance, but relatively poor integrated reasoning ability.
•  Type III: Ultra-high memorization and high integrated reasoning ability. This category includes GPT-4 and the three Chinese-oriented LLMs that exceed a size of 30B.
@@
In addition, we also evaluated the correlation between memorization and integrated reasoning during the LLM pre-training process, details can be found in Figure 10 in Appendix H.2.
@@ 16 LLaMA-2-13B LLaMA-2-70B (↓1) LLaMA-2-70B (↓1)
Table 7: Leaderboard on the MRI tasks. We propose two methods, i.e. FRMM and MIC, to compare the LLMs’ memorization-independent reasoning, as detailed in Appendix I. The arrows and numbers in brackets in the last two columns indicate changes in ranking order relative to the second column.
Methods We proposed two methods, FRMM and MIB, to compare the LLMs’ memorization-independent reasoning on the MRI tasks. The results are in Table 7.
•  Filtering Reasoning questions based on Mono-LLM-Memorization (FRMM) For each LLM, we selected reasoning questions based on its performance in memorization tasks: only retaining reasoning questions for which all related memorization questions were answered correctly. Then we calculated the accuracy of the retained reasoning questions for each LLM. The LLMs were then ranked based on the accuracy, producing the leaderboard shown in the penultimate column of Table 7. The detail of the FRMM is in Appendix I.1.
•  Memorization-Independent Battles among LLMs (MIB) Inspired by the pairwise battle method adopted in LLM evaluation Zheng et al. (2023), we tallyed each LLM’s performance in a “round-robin” tournament of pairwise match-ups and then ranked the performance of the LLMs. Specifically, we selected two LLMs at each time and filter the MRI task’s reasoning questions based on the performance of these two LLMs in memorization tasks. We only retained the reasoning questions whose related memorization questions were correctly answered by both LLMs. In this way, the two LLMs were battled under fair conditions. Then we calculated the accuracy of the two LLMs on the retained reasoning questions, and computed the difference in accuracy as the battle score between the two LLMs. For a total of 19 LLMs, we averaged each LLM’s scores from the 18 battles they participated in as their final scores. The LLMs were then ranked based on these final scores, producing the leaderboard shown in the last column of Table 7. The detail of the MIB is in Appendix I.2.
Analysis As shown in Table 7, when comparing the leaderboards for integrated and memorization-independent reasoning, Type III LLMs rank at the forefront and the Type I rank at the end in all leaderboards. There is a slight variation in the ranking order within the three types of LLMs.
Error Types For the in-depth analysis, we chose Vicuna-13B, Qwen-7B, Qwen-72B as representatives for Type I, II, and III LLMs, and filtered out the reasoning questions in the MRI tasks, only keeping those with correct answers to the related memorization questions, same as the FRMM in Appendix I.1. This ensured the LLM had sufficiently memorized the commonsense knowledge required for the retained reasoning questions, thereby minimizing the impact of memorization on reasoning. There are totally 500 reasoning questions in the 4 MRI tasks, and the numbers of the retained are 106, 323 and 402 for Vicuna-13B, Qwen-7B and Qwen-72B respectively, as shown in Table 8.
If the LLMs provided incorrect answers for the retained reasoning questions, these errors can be referred to as memorization-independent reasoning errors. We conducted the manual review and analysis of their reasoning process, and classified the errors into 4 main categories.
@@
We show examples of each type of errors in Figure 12 in Appendix J. The distribution of these error types are shown in Figure 4.
@@
Vicuna-13B-v1.5 Type I 500 106 54
Qwen-7B Type II 500 323 117
Qwen-72B Type III 500 402 63
@@ 6 Conclusion
This paper introduces CHARM, the first benchmark designed to comprehensively and thoroughly evaluate LLMs’ commonsense reasoning in Chinese. CHARM encompasses two counterpart commonsense domains, global and Chinese-specific, with the carefully selected tasks. We evaluated the representative prompt strategies for improving LLMs’ reasoning ability, and the empirical findings significantly enhances and supplements the conclusions of previous studies. CHARM comprises closely-interconnected reasoning and memorization tasks, helping to reveal the intrinsic correlation between memorization and reasoning of LLMs. We evaluated the strengths and weaknesses of different LLMs and conducted the detailed analysis of memorization-independent reasoning abilities. We hope that CHARM’s approach to studying the correlations between memoriztion and reasoning can serve as a reference for similar research in other fields.
@@
The automation of the construction and evaluation of CHARM needs further improvement, including the following: (1) Most of the questions in CHARM Chinese domain are manually constructed by the author. This limits the number of benchmark questions and the range of knowledge pieces covered. (2) Regarding memorization-independent reasoning, we chose only 3 LLMs as representative and manually categorized the types of errors within CHARM. In future research, we could employ robust LLMs, like GPT-4, for automated error classification and statistical analysis.
@@
This work involved human annotation. We have provided appropriate compensation for all annotators. The total cost of annotation for the project is about 2.2k RMB. For all annotators, we explicitly informed them about the use of the data and required them to ensure that the questions included in CHARM do not involve any social bias, ethical issues or privacy concerns during the annotation process.
Acknowledgements
We thank Dr. Qipeng Guo of Shanghai AI Lab for his invaluable suggestions. We additionally thank our crowd annotators for their diligent work and the anonymous reviewers for their constructive comments.
@@ Appendix C Question Examples of the Memorization Task in CHARM
Appendix D Details of Quality Assurance
D.1 Quality Assurance Procedures
After construction of CHARM, we conducted the quality assurance to ensure the quality of the questions. We hired professional NLP task annotators to review the questions we have constructed. The entire quality assurance includes the following steps:
Step(1) Annotation Task Submission and Assignment: We packaged the constructed questions into annotation task packages. Usually, one annotation package corresponds to all the questions of one reasoning or memorization task in Table 2. Before submitting the annotation tasks, we had written the annotation requirement document, which includes key requirements and typical examples. After the annotation tasks were submitted to the online annotation platform, they were assigned to suitable annotators. Typically, a task would be assigned to either two or four annotators.
Step(2) Annotator training: After the annotation task was assigned, we organized an online meeting to train the annotators. We specifically asked annotators to avoid social bias and sensitive issues such as morality.
Step(3) Trial review: Before the official review begun, we randomly selected 20% of the questions for the annotators to try annotating. We reviewed the results of the trial annotation, corrected any detailed issues or understanding deviations in the review process in a timely manner, to ensure the quality of subsequent reviews.
Step(4) Official review: For each question, whether handcrafted or translated, we had two annotators do independent reviews. We only provided the question to the annotators, not our answers. Using external resources fully (such as search engines, online encyclopedias, etc.), the annotators would provide the answers. Annotators could also provide feedback on issues with the question itself (for example, translation errors or questions that do not meet the commonsense standards). Only when both annotators believed that there was no problem with the question, and their provided answers were consistent with our previous answers, did we consider the question to have passed the review.
Step(5) Consultation and correction: For questions that did not pass the step(4) review, several authors would hold a meeting to discuss the questions. There are three cases in total:
•  Case(1) Retain: There was no problem with the question itself, but the annotator answered incorrectly. We retained these questions.
•  Case(2) Discard: The question had significant errors due to translation issues or problems with the question itself. We discarded these questions.
•  Case(3) Correct: There were minor issues with the question or the answer, the multiple authors would discuss together and complete the correction of the questions.
D.2 Information of the Annotators
We submitted the annotation task online to the professional data annotation company, which organized the annotators to complete the annotation work. A total of 30 professional annotators, all native Chinese speakers with extensive experience in natural language processing tasks, were involved in this project. They possess the expertise to discern and comprehend commonsense knowledge pertinent to both global and Chinese-specific contexts. Here is the specific information about these annotators.
Education and profession: Of the 30 annotators, 11 have a bachelor’s degree and 19 have an associate degree. Regarding their fields of study, 9 are in humanities (4 specialized in design, 3 in business, 2 in language), while 21 are in science and engineering (11 specialized in computer science, 4 in automation, 3 in medicine, 2 in architecture, 1 in mathematics).
Age: 14 annotators are aged 20-25, 8 are aged 26-30, 7 are aged 30-35, and 1 is aged 35-40.
The annotators are all located in Changzhi City, Shanxi Province, China. We offer the hourly wage of 23.75 RMB for each annotator, which is higher than the local minimum wage standard.
D.3 Other Details
During the construction of the CHARM project, we submitted a total of 22 annotation tasks, which together contained approximately 3.55k questions, and took a total of approximately 93 hours of the annotators’ working time. The entire process for each annotation task, from step(1) to step(4), typically required half a day. It needs to be emphasized that the annotators were only tasked with checking and answering questions that we have already created; they were NOT responsible for creating questions from scratch.
The annotation work for the entire project spanned from January 2 to January 26, 2024. Each annotation task required only a few annotators to complete, rather than all 30 annotators. During the project period, the annotators participating in this project were also undertaking other data annotation tasks in the companies. In fact, our annotation tasks only accounted for a very small part of their working time.
The statistics results of the quality evaluation process are as followings:
•  For reasoning tasks, the average failure rate in step(4) is 0.19, and the average ratio of case(3) in step(5) is 0.04.
•  For memorization tasks, the average failure rate in step(4) is 0.04, and the average ratio of case(3) in step(5) is 0.02.
Appendix E Examples of Prompt Strategies
@@
Appendix F Detailed Evaluation Results of 19 LLMs with 5 Prompt Strageties on Reasoning Tasks
@@
Appendix G Performance of LLMs on Chinese Commonsense Knowledge Aspects
@@
Appendix H Correlation of Memorization and Integrated Reasoning
H.1 Detailed Correlations of Memorization and Integrated Reasoning on the 4 MRI Tasks
@@
H.2 Correlation of Memorization and Integrated Reasoning throughout the LLM pretraining
@@
With the increase in the number of tokens during the training process, the model’s memorization ability quickly reach a high level (in fact, there is no particularly obvious difference between the results of the first checkpoint and the final results). This is because the knowledge involved in our task setting is the most basic commonsense, and thus widely and abundantly exists in various Chinese training corpora.
@@
Appendix I Leaderboard of Memorization-Independent Reasoning
It is non-trivial to acquire and compare the memorization-independent reasoning abilities of the LLMs. Intuitively, we can filter the reasoning questions by only retaining those whose related memorization questions are all correctedly answered by every LLMs. This approach ensures that each LLM has memorized the commonsense knowledge necessary for the retained reasoning questions. However, when we applied this process to all the 19 LLMs, only 28 reasoning questions remained out of the original 500 in the MRI tasks, which was obviously insufficient in number and lacks diversity, thereby introducing a high degree of uncertainty due to randomness.
Therefore, we proposed two slightly more complex methods, one called Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM), the other is Memorization-Independent Battles among LLMs (MIB).
I.1 Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM)
This method is relatively simple, but has some flaws to a certain extent. For each LLM, we selected reasoning questions based on its performance in memorization tasks: only retaining reasoning questions for which all related memorization questions are answered correctly. It’s clear that, after individual filtration, different LLMs would retain different reasoning questions, and even differ in the number of retained reasoning questions, as shown in the “# retained” column in Table 12.
Then, we calculated the accuracy of the retained reasoning questions for each LLM, and the results are shown in the “Retained Acc” column in Table 12. The LLMs were then ranked based on the accuracy, producing the leaderboard shown in the penultimate column of Table 7.
@@ I.1 Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM)
To overcome this, we proposed the MIB method.
I.2 Memorization-Independent Battles among LLMs (MIB)
To overcome the shortcomings of the FRMM method, we referred to the pairwise battle method adopted in LLM evaluation Zheng et al. (2023). By tallying each LLM’s performance in a “round-robin” tournament of pairwise match-ups, we ranked the performance of the LLMs.
Specifically, we selected two LLMs at each time and filtered the MRI task’s reasoning questions based on the performance of these two LLMs in memorization tasks. We only retained the reasoning questions whose related memorization questions were correctly answered by both LLMs. In this way, the two LLMs are battled under fair conditions. We then calculated the accuracy of these two LLMs on the retained reasoning questions separately, and compute the difference in accuracy as the battle score between the two models.
As shown in Figure 11, the element $E_{ij}$ represents the accuracy of LLM $i$ minus the accuracy of LLM $j$ during the battle between the two LLMs. For a total of 19 LLMs, we averaged each LLM’s scores from the 18 battles they participated in as their final scores, as shown in Table 13.
Finally, we ranked the LLMs based on these scores to produce the leaderboard shown in the last column of Table 7.
Appendix J Memorization-Independent Reasoning Errors